### v1

@@ Appendix C Proofs for Section 4
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0},W^{\prime})$, and define
$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}.$
We verify that $W^{*}$ is a regular subpopulation of $W^{\prime}$. First, we see that $\mathbb{P}(W^{*}=1)>0$ because
$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)+\mathbb{P}(W^{*}=1\mid W^{\prime}=0)\mathbb{P}(W^{\prime}=0)$
$=\mathbb{P}\left(\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}=1\mid W^{\prime}=1\right)\mathbb{P}(W^{\prime}=1)$
$=\mathbb{E}\left[\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
$=\mathbb{E}\left[\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
$=\frac{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
$=\frac{\mathbb{E}[a(X)\mid W_{0}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
$>0.$
The second equality follows from $\mathbb{P}(W^{*}=1\mid W^{\prime}=0)=0$, which holds by the construction of $W^{*}$. The fourth equality holds from $U\perp\!\!\!\perp(X,W^{\prime})$ and from $\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\in[0,1]$ almost surely given $W^{\prime}=1$. To establish the final inequality, recall that $\mathbb{P}(W^{\prime}=1)$ and $\mathbb{E}[a(X)\mid W_{0}=1]$ are positive by assumption. Also $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))\leq\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ since $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1)$ is a subset of $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1)$. That $W^{*}$ satisfies the two properties of Definition 3.1 holds immediately. Therefore, $W^{*}$ is a regular subpopulation of $W^{\prime}$.
Finally, let
$\underline{w}^{*}(X)$ $=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$
$=\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)$
$=\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}.$
Letting $\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]=\tau_{0}(X)\in\mathcal{T}_{\text{all}}$ and using Proposition 3.1, we can see that
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{\prime}(X)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
$=\mu(a,\tau_{0}).$
Since $\tau_{0}\in\mathcal{T}_{\text{all}}$ was arbitrary, we have that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ which concludes the proof.
Part 2: $\mathcal{T}=\{\tau_{0}\}$
To simplify the notation in the proof, we let $\mu\equiv\mu(a,\tau_{0})$.
()($\Longrightarrow$)( ⟹ ) First, let $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. Since $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)=0$, which implies $\mathbb{P}(\tau_{0}(X)>\mu\mid W^{\prime}=1)=1$. Since $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$, we can write by Proposition 3.1
$\mu$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
$\geq\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mu\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}=\mu.$ (C.1)
The inequality is strict unless $\mathbb{E}[(\underbrace{\tau_{0}(X)-\mu}_{>0\text{ w.p.1}})\underbrace{\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)}_{\in[0,1]}\mid W^{\prime}=1]=0$ holds. This holds if $\mathbb{P}((\tau_{0}(X)-\mu)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$, which in turns occurs if and only if $\mathbb{P}(\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$. This implies $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)=0$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)=0$, a contradiction of $W^{*}\in\mathcal{W}(a,W^{\prime},\{\tau_{0}\})$. Therefore, the inequality in (C.1) is strict and yields $\mu>\mu$, a contradiction. Therefore, $\mathcal{W}(a;W^{\prime},\{\tau_{0}\})=\emptyset$ when $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$.
@@
$\mathcal{X}^{-}$ =$=\{x\in\text{supp}(X):\tau_{0}(x)\leq\mu\}\qquad\text{ and }\qquad\mathcal{X}^{+}=\{x\in\text{supp}(X):\tau_{0}(x)\geq\mu\}.$
By $\mu\in\mathcal{S}(\tau_{0};W^{\prime})$, $\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)>0$. Similarly, $\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)>0$.
Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W^{\prime},W_{0})$. For $u\in[0,1]$ let
$W^{*}(u)$ $=(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+}))\cdot W^{\prime}.$
We can see that $W^{*}(u)\in\{0,1\}$, that $W^{*}(u)\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$, and that $\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=0)=0$. To show that $W^{*}(u)$ characterizes a regular subpopulation of $W^{\prime}$, we also show that it is nonzero with positive probability:
$\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)$ $=\mathbb{P}(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+})=1\mid W^{\prime}=1)$
$=\mathbb{P}(U>u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)+\mathbb{P}(U\leq u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)$
$=(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)$
$>0$
for all $u\in[0,1]$. Therefore, $\mathbb{P}(W^{*}(u)=1)=\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)>0$ by $\mathbb{P}(W^{\prime}=1)>0$. Hence, $W^{*}(u)\in\text{SP}(W^{\prime})$ for all $u\in[0,1]$.
For a given $u$, we have that $\underline{w}^{*}(X;u)\equiv\mathbb{P}(W^{*}(u)=1\mid X,W^{\prime}=1)=(1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+})$. Therefore, using Proposition 3.1,
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X;u)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{*}(X;u)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\mid W^{\prime}=1]}$
$=\frac{(1-u)\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]+u\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}.$
By construction, $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{-})\leq\mu\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{+})\geq\mu\mathbbm{1}(X\in\mathcal{X}^{+})$ almost surely. Therefore,
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(0)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}\leq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}=\mu$
and
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(1)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}\geq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}=\mu.$
By the continuity of $\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ in $u$ and the intermediate value theorem, there exists $u^{*}\in[0,1]$ such that $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u^{*})=1]$ and $W^{*}(u^{*})\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. ∎
Proof of Theorem 4.4.
Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
First suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$. From Theorem 4.3, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Written differently, we have that
$\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$ $=\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1\right]}$
for all $\tau_{0}\in\mathcal{T}_{\text{all}}$. From derivations in the proof of Theorem 4.3, we have that $\mathbb{P}(C\cdot\frac{a(X)}{\underline{w}^{\prime}(X)}=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1)=1$ for some positive constant $C>0$.
Since $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\leq 1$ almost surely given $W^{\prime}=1$, we must have $C\cdot a(X)/\underline{w}^{\prime}(X)\leq 1$ almost surely given $W^{\prime}=1$. This means $C$ is bounded above by $\inf(\text{supp}(\underline{w}^{\prime}(X)/a(X)\mid W^{\prime}=1))$, which is strictly positive by assumption. Therefore,
$\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ $=\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$
$\leq\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\right]}{\mathbb{P}(W^{\prime}=1)}$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)$
$\qquad\cdot\mathbb{E}\left[\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid X,W_{0}=1\right]\mid W_{0}=1\right]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\cdot\mathbb{E}[a(X)\mid W_{0}=1]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}.$
The fifth line follows from $W^{\prime}$ being a subpopulation of $W_{0}$, and the last line follows from the law of iterated expectations.
This upper bound is sharp because it is attained by setting
$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}$
and noting that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ from Theorem 4.3.
Now suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$. By Theorem 4.3, $\mathcal{W}(a;W^{\prime},\mathcal{T})=\emptyset$ and therefore $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})$ is zero.
Part 2: $\mathcal{T}=\{\tau_{0}\}$
In this case, we seek to maximize $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ subject to $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. Using Proposition 3.1, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ as $\mathbb{E}[\tau_{0}(X)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]/\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$. The result then follows from a direct application of Theorem 4.2 that replaces $W_{0}$ by $W^{\prime}$ in its statement and $\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ by $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$ in the proofs. ∎
Appendix D Difference-in-Differences

### v3

@@ Appendix G Proofs for Appendix C
by the continuity of $\psi$ and Assumption C.1. ∎
Proof of Theorem C.2.
@@
Let $h=(h_{1},h_{2},h_{3})$ and $h^{\prime}=(h_{1}^{\prime},h_{2}^{\prime},h_{3}^{\prime})$.
$|\widehat{\psi}(h^{\prime})-\widehat{\psi}(h)|$ $\leq\left|\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}(h_{1}^{\prime}(j)-h_{1}(j))\right|$
$+\left|\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\left(\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}^{\prime}(j)-\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)\right)\right|$
$+\left|\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}(h_{2}^{\prime}(j)-h_{2}(j))\right|$
$+\left|\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{w}_{0}(x_{j})}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}(h_{3}^{\prime}(j)-h_{3}(j))\right|$
$\leq\left(\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})^{2}\widehat{p}_{j}^{2}}{\widehat{\mathbb{P}}(W_{0}=1)^{2}\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\right)^{1/2}\|h_{1}^{\prime}-h_{1}\|$ (A.6)
$+\frac{|\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]|}{\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\left|\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}^{\prime}(j)-\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)\right|$ (A.7)
$+\left(\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])^{2}\widehat{p}_{j}^{2}}{\widehat{\mathbb{P}}(W_{0}=1)^{2}\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\right)^{1/2}\|h_{2}^{\prime}-h_{2}\|$ (A.8)
$+\left(\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])^{2}\widehat{w}_{0}(x_{j})^{2}}{\widehat{\mathbb{P}}(W_{0}=1)^{2}\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\right)^{1/2}\|h_{3}^{\prime}-h_{3}\|,$ (A.9)
where we applied the Cauchy–Schwarz inequality several times. Note that the maximum function is Lipschitz with Lipschitz constant one and therefore
$\left|\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}^{\prime}(j)-\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)\right|$ $\leq\sum_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}|h_{1}^{\prime}(j)-h_{1}(j)|$
$\leq\sum_{j=1}^{K}|h_{1}^{\prime}(j)-h_{1}(j)|$
$\leq\sqrt{K}\|h_{1}^{\prime}-h_{1}\|.$
[…]
Thus we now consider the consistency of the different components of $\widehat{\psi}(h)$. Applying Theorem C.1 and the continuous mapping theorem, we can show that terms
$\left(\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}h_{1}(j),\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}},\right.$
$\left.\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}h_{2}(j),\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{w}_{0}(x_{j})}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}h_{3}(j)\right),$
are all consistent for their counterparts in $\psi(h)$. It remains to show that $\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)=\max_{j\in\Psi_{\mathcal{X}^{+}}}h_{1}(j)+o_{p}(1)$. This holds if the set $\widehat{\Psi}_{\mathcal{X}^{+}}$ is consistent for $\Psi_{\mathcal{X}^{+}}$, which we establish here. Let $k\in\Psi_{\mathcal{X}^{+}}$. Then,
$\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})$ $=\mathbb{P}\left(\widehat{a}(x_{k})\geq\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j})-\xi_{n}\right)$
$=\mathbb{P}\left(\sqrt{n}(\widehat{a}(x_{k})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))\geq-\sqrt{n}\xi_{n}\right)$
$=\mathbb{P}\left(\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))\geq-\sqrt{n}\xi_{n}\right).$
The third equality follows from $k\in\Psi_{\mathcal{X}^{+}}$. By the proof of Theorem C.1, $\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))=o_{p}(1)$. Since $-\sqrt{n}\xi_{n}\rightarrow-\infty$, $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})\to\mathbb{P}(0\geq-\infty)=1$ when $k\in\Psi_{\mathcal{X}^{+}}$.
Now suppose that $k\notin\Psi_{\mathcal{X}^{+}}$. Then,
$\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})$ $=\mathbb{P}(\widehat{a}(x_{k})\geq\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j})-\xi_{n})\rightarrow\mathbb{P}(a(x_{k})\geq\max_{j\in\mathcal{X}^{+}}a(x_{j})-0)=0,$
[…]
Appendix H Additional Derivations for Difference-in-Differences